OpenStudy (itz_sid):

HELP PLEASE! D:

OpenStudy (itz_sid):

I need help with #5.

OpenStudy (itz_sid):

Here is my work.

OpenStudy (itz_sid):

Idk, it looks right to me. D:

OpenStudy (itz_sid):

Hm?

OpenStudy (irishboy123):

oooops it's shiτ

OpenStudy (itz_sid):

???

OpenStudy (irishboy123):

well, if it takes 26N to stretch it by 6cm, then \(k = \dfrac{26}{0.06} = \dfrac{13}{0.03} \frac{N}{m}\) The work to stretch it by 3cm is then equal to the energy stored in the spring: \(= \dfrac{1}{2} k x^2 = \dfrac{1}{2} \dfrac{13}{0.03} (0.03)^2\) that's pretty much all there is to it

OpenStudy (itz_sid):

I got 0.19 as my answer. And got it wrong :/

OpenStudy (irishboy123):

type it into your calculator again

OpenStudy (irishboy123):

Or nvm :-))

OpenStudy (itz_sid):

Ugh, Now i only have 1 attempt left at the right answer. Can someone please help? :(

OpenStudy (loser66):

Which course are you in? calculus 2 or differential equation?

OpenStudy (itz_sid):

Calculus 2

OpenStudy (loser66):

ok. Just want to know!! :)

OpenStudy (itz_sid):

Do you know how I can solve this?

OpenStudy (loser66):

I don't promise anything, but will try all my best.

OpenStudy (irishboy123):

no pressure

OpenStudy (loser66):

to me, k is what Irishboy said, just 26/0.06 and f(x) is just kx, and the limit is from 0 to 0.03.

OpenStudy (irishboy123):

pas de pressure

OpenStudy (loser66):

oh, it says you have to have it upto 2 decimal points, so, your answer should be 0.20

OpenStudy (irishboy123):

LOL!!

OpenStudy (loser66):

I saw a lot of cases like this. The computer is so sensitive. I don't know!! :)

OpenStudy (itz_sid):

Wow...... lmao. Thanks xD

OpenStudy (loser66):

did you get it right?